# Reciprocity-Aware Convolutional Neural Networks for Map-Based Path Loss Prediction

### v1

@@ I Introduction
Modern path loss models rely on empirical, data-driven approaches to achieve low root mean squared error (RMSE) during inference in unseen geographic regions. With enough data, neural networks can be efficient for modeling complex, non-linear relationships [2, 3]. While data-driven approaches can be helpful for such relationships, the resulting models can only be as powerful as their training data. More diverse training data can often help neural networks reduce RMSE in both training and inference [3].
@@
Figure 1: All input channels (re-sampled and normalized) visualized for a 411 m link. (a) Direct path. (b) Distance from Tx to each pixel. (c) Surface. (d) Frequency (449 MHz, normalized to 0.056125).
@@
This paper investigates the use of data augmentation for training map-based path loss models that are aware of reciprocity, providing low RMSE for multiple link scenarios. Section II describes two measurement datasets used to train and evaluate models in multiple link scenarios. Section III begins with an overview of the transformations applied to represent various scenarios, followed by testing on multiple scenarios not seen in training. Section IV describes the use of these transformations to augment the training, and is followed by Section V, which discusses the results. Section VI concludes the paper.
@@ II Measurement Data
Holding out entire drive tests is a form of geographic cross-validation [10], providing a proper assessment of model generalization to unseen regions. Geographic cross-validation is necessary for this data since different regions can have varied statistics in their DSM data, and an individual region’s DSM data can often identifiable characteristics that are common throughout the region. Therefore, training and testing within the same region(s) can provide an optimistic assessment of performance due to spatial autocorrelation [11]. Holding out complete drive tests ensures that the assessment is valid by preventing data leakage between adjacent samples from training to test within a region. More information on the data cleaning, cross-validation approach, and hyperparameters can be found in [4].
In this paper, we also use 3455 MHz measurement data taken by Communications Research Centre Canada (CRC) across 15 measurement sites in Ottawa [12] to evaluate models on BS-to-BS scenarios. Tx and Rx heights in this dataset are both 11 m above ground. The dataset comprises 71 measurements after completing the filtering steps described in [4].
Table I summarizes both measurement datasets and their usage in the paper.
@@
Measurement filtering Above system noise floor by 6 dB margin Below maximum measurable path loss
@@
*For all drive tests except London, which used a 25 m Tx.
**Measurement data is also available for Scar Hill, but high-resolution DSM data was not available at the time of publication.
As per guidelines to reduce systematic measurement errors to 1 dB or lower [13].
@@ III Data Representations
The Ofcom data provides BS-to-UE measurements. These measurements will be used to represent multiple link scenarios in both training and evaluation. Conversely, the CRC data provides BS-to-BS measurements. These will be used for evaluation.
@@ III-A Ofcom Data Transformations
The Ofcom data provides BS-to-UE measurements. In this paper, we transform the data to represent the opposite scenario: UE-to-BS. To this end, we denote the following two transformations:
@@ III-B CRC Measurement Data
The CRC measurement data provides measurements of BS-to-BS scenarios, where both Tx and Rx are antennas fixed at a tall height. Fig. 4 shows an example of a re-sampled and normalized path profile derived from the CRC data, for a 1 km link. Table II summarizes the dataset representations for all path profiles.
@@
Figure 4: Ottawa path profile visualized using re-sampled and normalized CNN inputs. (a) Surface channel visualized in 2-D. (b) All channels visualized in 1-D.
@@ III-C Model Evaluation on Multiple Test Scenarios
Reflection is conducted on the test sets (holdouts) from the original cross-validation from [4]. The corresponding trained CNN is evaluated on each of these new transformed test sets (UE-to-BS), and compared with the original results (“identity” transformation, BS-to-UE) in Table III. Each CNN is also evaluated on the BS-to-BS samples from Ottawa [12].
@@ IV Training Set Augmentation
The model is re-trained with augmented training/validation sets. $n$ samples from each of the five training/validation cities are duplicated and reflected for multiple values of $n$ (resulting in 12 000 + $n$ samples per city between training and validation). This is employed for varying values of $n$ (no augmentation, 4%, and 25%). As described in [4], each cross-validation is conducted 10 times. The holdouts use 12 000 samples from the held out drive test for both the reflected and original test sets. The resulting models are evaluated on both the original and reflected test sets, as well as the CRC data. The results are shown in Tables IVV, and VI respectively. The results are summarized in Table VII. Finally, Fig. 5 visualizes the error distribution across all holdouts using kernel density estimation.
@@
Figure 5: Kernel density estimation of prediction bias. (a) Identity (BS-to-UE). (b) Reflection (UE-to-BS). (c) CRC data (BS-to-BS).
@@ V Discussion
The data augmentation has had a significant effect on test RMSE. Adding just 500 simulated uplink samples to each city (just 4% of the original training/validation set) reduces uplink RMSE from 16.20 dB to 7.76 dB. Even the mean RMSE on the unaltered (downlink) test set is reduced from 7.35 dB to to 7.32 dB, likely benefiting from the increased sample diversity. Mean RMSE on the CRC (BS-to-BS) test set is also improved with training augmentation, from 7.33 dB to 7.16 dB on average. Increasing the number of transformed training/validation samples to 3000 per city (25% of the original data) incrementally improves the performance on the transformed test set, to 7.42 dB mean RMSE. While the RMSE on the unaltered holdouts stays roughly the same (within 0.01 dB), RMSE on the BS-to-BS test set is improved to 7.09 dB.
Fig. 5 demonstrates that without training representation, path loss is underestimated in uplink scenarios. One potential explanation is that the majority of obstructions are near the Rx in the downlink scenarios. The Tx is often above the nearby buildings while the Rx is surrounded by buildings and trees. This type of scenario being over-represented in training may have caused the models to underemphasize the area around the Tx while focusing on the area around the Rx. Consequently, the model may have overlooked the obstructions near the Rx in the uplink samples, resulting in a much lower predicted path loss than reality. Augmenting 4% of training samples appears to have resolved the issue by allowing the model to learn that obstructions near both Tx and Rx can impact path loss. The $n_{1}$ models have a mean error of 0.62 dB on the reflected test samples – a significant improvement over --12.31 dB.
@@
The results discussed thus far showcase the effect of training data augmentation on generalization, by comparing RMSE. However, the original purpose of this study was to examine whether the CNNs can accurately model reciprocity. Therefore, it is beneficial to examine the individual predictions to see how well they align between downlink and uplink scenarios. Fig. 6 shows the distribution of the difference between the identity prediction and the reflected prediction for each test link. For an ideal model with perfect reciprocity, the distribution would be a unit impulse at $x=0$, indicating that each sample has the same prediction in both directions.
Figure 6: Kernel density estimation of prediction on downlink minus prediction on uplink.
The original models have a high mean and standard deviation, showing both a lack of consistency and also a tendency to predict higher path loss in BS-to-UE samples than the reverse, which is consistent with the fact that path loss is under-predicted in UE-to-BS samples. The first training data augmentation decreases the mean difference from 12.50 dB to --0.05 dB, and the SD from 9.48 dB to 4.74 dB, reinforcing the power of training representation. The second training data augmentation has a more subtle yet still relevant improvement, to a mean and SD of --0.02 dB and 3.90 dB, respectively. The improvements provide diminishing returns, but still show an incremental improvement by tightening the distribution as individual uplink and downlink predictions converge.
@@ VI Conclusion
Future work should investigate the discussed techniques in a wider variety of scenarios, including vertical use cases, such as vehicular communications. Additionally, this approach should be validated in inference using uplink measurements. This would provide increased confidence on the model’s ability to handle varied communications scenarios, which may not always be truly reciprocal in practice. A dedicated study may also determine the optimal number of augmented samples to balance test RMSE on multiple link scenarios. Finally, future research could also explore explicitly enforcing reciprocity constraints via alternative model architectures and/or loss functions, and compare the effectiveness against the data augmentation methods presented here.

### v2

@@ I Introduction
Modern path loss models rely on empirical, data-driven approaches to achieve low root mean squared error (RMSE) during inference in unseen geographic regions. With enough data, neural networks can be efficient for modeling complex, nonlinear relationships [2, 3]. While data-driven approaches can be helpful for such relationships, the resulting models can only be as powerful as their training data. More diverse training data can often help neural networks reduce RMSE in both training and inference [3].
@@
Figure 1: All input channels (re-sampled and normalized) visualized for a 411 m link. (a) Direct path. (b) Distance from Tx to each pixel. (c) Surface. (d) Frequency. Channels are separately normalized for visualization.
@@
This paper investigates the use of data augmentation for training map-based path loss models that are aware of reciprocity, providing low RMSE for multiple link scenarios. Unlike [4] and [6] which are limited to downlink scenarios, this paper evaluates models in more diverse links. Section II describes two measurement datasets used to train and evaluate models in various link scenarios. Section III begins with an overview of the transformations applied to represent various scenarios, followed by testing on multiple scenarios not seen in training. Section IV describes the use of these transformations to augment the training, and is followed by Section V, which discusses the results. Section VI concludes the paper.
@@ II Measurement Data
Holding out entire drive tests is a form of geographic cross-validation [10], providing a proper assessment of model generalization to unseen regions. Geographic cross-validation is necessary for this data since different regions can have varied statistics in their DSM data, and an individual region’s DSM data can often identifiable characteristics that are common throughout the region. Therefore, training and testing within the same region(s) can provide an optimistic assessment of performance due to spatial autocorrelation [11]. Holding out the drive tests ensures rigorous assessment by preventing data leakage between adjacent samples from training to test within a region. More information on data cleaning, cross-validation approach, and hyperparameters can be found in [4].
In this paper, we also use 3455 MHz measurements taken by Communications Research Centre Canada (CRC) across 15 measurement sites in Ottawa [12] to evaluate models on BS-to-BS scenarios. Tx and Rx heights in this dataset are both 11 m above ground. The dataset comprises 71 measurements after completing the filtering steps described in [4]. Table I summarizes both measurement datasets.
@@
Measurement filtering Above system noise floor by 6 dB margin** Below maximum measurable path loss
@@
*Measurement data is also available for Scar Hill, but high-resolution DSM data was not available at the time of publication.
**As per guidelines to reduce systematic measurement errors to 1 dB or lower [13].
For all drive tests except London, which used a 25 m Tx.
@@ III Data Representations
The Ofcom data provides BS-to-UE measurements. These measurements will be used to represent multiple link scenarios in both training and evaluation. Conversely, the CRC data provides BS-to-BS measurements, to be used for evaluation.
@@ III-A Ofcom Data Transformations
The Ofcom data provides BS-to-UE measurements. In this paper, we transform the data to represent the opposite scenario: UE-to-BS. To this end, we define the following two transformations:
@@ III-B CRC Measurement Data
The CRC dataset provides measurements of BS-to-BS scenarios, where both Tx and Rx are antennas fixed at a tall height. Fig. 4 shows a re-sampled and normalized path profile derived from the CRC data, for a 1 km link. Table II summarizes the dataset representations for all path profiles.
@@
Figure 4: Ottawa path profile visualized using re-sampled and normalized CNN inputs. (a) Surface visualized in 2-D. (b) All channels visualized in 1-D.
@@ III-C Model Evaluation on Multiple Test Scenarios
Reflection is conducted on the test sets (holdouts) from the original cross-validation from [4]. The corresponding trained CNN is evaluated on each of these new transformed test sets (UE-to-BS), and compared with the original results (“identity” transformation, BS-to-UE), via mean and standard deviation (SD) in Table III. Each CNN is also evaluated on the BS-to-BS samples from Ottawa [12].
@@ IV Training Set Augmentation
The model is re-trained with augmented training/validation sets. $n$ samples from each of the five training/validation cities are duplicated and reflected for multiple values of $n$ (resulting in 12 000 + $n$ samples per city between training and validation). This is employed for varying values of $n$ (no augmentation, 4%, and 25%). As described in [4], each cross-validation is conducted 10 times. The holdouts use 12 000 samples from the held out drive test for both the reflected and original test sets. The resulting models are evaluated on both the original and reflected test sets, as well as the CRC data. The results are shown in Tables IVV, and VI respectively. The results are summarized in Table VII. Finally, Fig. 5 visualizes the error distribution across all holdouts using kernel density estimation (KDE).
@@
Figure 5: KDE of prediction bias. (a) Identity (BS-to-UE). (b) Reflection (UE-to-BS). (c) CRC data (BS-to-BS).
@@ V Discussion
The data augmentation has had a significant effect on test RMSE. Adding just 500 simulated uplink samples to each city (just 4% of the original training/validation set) reduces uplink RMSE from 16.20 dB to 7.76 dB. Even the mean RMSE on the unaltered (downlink) test set is reduced from 7.35 dB to to 7.32 dB, likely benefiting from the increased sample diversity. Mean RMSE on the CRC (BS-to-BS) test set is also improved with training augmentation, from 7.33 dB to 7.16 dB. Increasing the number of transformed training/validation samples to 3000 per city (25% of the original data) incrementally improves the performance on the transformed test set, to 7.42 dB mean RMSE. While the RMSE on the unaltered holdouts stays roughly the same as $n_{0}$ (within 0.01 dB), RMSE on the BS-to-BS test set is improved to 7.09 dB.
Fig. 5 demonstrates that without training representation, path loss is underestimated in uplink scenarios. One potential explanation is that the majority of obstructions are near the Rx in the downlink scenarios. The Tx is often above the nearby buildings while the Rx is surrounded by buildings and trees. This type of scenario being over-represented in training may have caused the models to underemphasize the area around the Tx while focusing on the area around the Rx. Consequently, the model may have overlooked the obstructions near the Rx in the uplink samples, resulting in a much lower predicted path loss than reality. Augmenting 4% of training samples appears to have resolved the issue by allowing the model to learn that obstructions near both Tx and Rx can impact path loss. The $n_{1}$ models have a mean error of 0.62 dB on the reflected test samples—a significant improvement over --12.31 dB.
@@
The results discussed thus far showcase the effect of training data augmentation on generalization, by comparing test RMSE. However, the original purpose of this study was to examine whether the CNNs can accurately model reciprocity. Therefore, it is beneficial to examine the individual predictions to see how well they align between downlink and uplink scenarios. Fig. 6 shows the distribution of the difference between the identity prediction and the reflected prediction for each test link. For an ideal model with perfect reciprocity, the distribution would be a unit impulse at $x=0$, indicating that each sample has the same prediction in both directions.
Figure 6: KDE of prediction on downlink minus prediction on uplink.
The original models have a high mean and SD, showing both a lack of consistency and also a tendency to predict higher path loss in BS-to-UE samples than the reverse, which is consistent with the fact that path loss is under-predicted in UE-to-BS samples. The first training data augmentation decreases the mean difference from 12.50 dB to --0.05 dB, and the SD from 9.48 dB to 4.74 dB, reinforcing the power of training representation. The second training data augmentation has a more subtle yet still relevant improvement, to a mean and SD of --0.02 dB and 3.90 dB, respectively. The improvements provide diminishing returns, but still show an incremental improvement by tightening the distribution as individual uplink and downlink predictions converge.
@@ VI Conclusion
Future work should investigate the discussed techniques in a wider variety of scenarios, including vertical use cases, such as vehicular communications. Additionally, this approach should be validated in inference using uplink measurements. This would provide increased confidence on the model’s ability to handle varied communications scenarios, which may not always be truly reciprocal in practice. A dedicated study may also determine the optimal number of augmented samples to balance training times with test RMSE on multiple link configurations. Finally, future research could also explore explicitly enforcing reciprocity constraints via alternative model architectures and/or loss functions, and compare the effectiveness against the data augmentation methods presented here.